### v1

@@ 1 Introduction
One promising approach to mitigating the curse of dimensionality lies in exploiting separability, that is, the ability to approximate a global function as a sum of local contributions involving only a subset of variables. Separability properties can be intrinsic to the problem under study, as in systems where interactions decay with distance or graph connectivity, such as in physical systems governed by locality principles or networked systems with sparse interactions. At a computationa level, separability can be enforced by an suitable approximation ansatz, including separation of variables, ANOVA[22], and most recent tools such as tensor decompositions[5, 21]. Theoretical and computational frameworks for separable approximations offer significant reductions in complexity while retaining essential structural features of the underlying functions.
@@
The concept of decaying sensitivity is intrinsically linked to the off-diagonal decay property, a topic extensively explored in the field of numerical linear algebra. A foundational theoretical framework for understanding this behavior is introduced in [3], where decay bounds are derived for the entries of general matrix functions. This analysis is further advanced in [12], which examines decay properties in the context of optimal control, specifically for solutions to Lyapunov equations. Additionally, related investigations are presented in [18], where the sparsity structure of solutions to Riccati equations is analyzed, providing complementary insights into the decay and sparsity phenomena in matrix equations. Our particular focus of application are optimal control problems, for which such decaying sensitivity properties in time and space have been intensively studied in the past years, see, e.g., [27, 20, 11] for decaying sensitivity in time and [26, 30, 23, 9] for decaying sensitivity in space for problems governed by discrete-time dynamics, ODEs and PDEs. See also [25] for a related property for static nonlinear optimization problems. The key prerequisite for the occurrence of decaying sensitivity in space is that the optimal control problem consists of subsystems that are interconnected by a graph structure. For instance in multi-agent control systems or robotic swarms, each agent’s state and control variables are influenced predominantly by neighboring agents, leading to a natural decaying sensitivity structure. By leveraging this property, we develop computationally efficient methods for approximating optimal value functions through separable structures.
@@ 1.1 Contribution
The main purpose of this paper is to show that decaying sensitivity in space enables the construction of accurate seperable approximations. While our result is formulated for arbitrary functions and thus provides a general framework for separable approximations, we focus on the approximation of value functions arising in the context of dynamic programming or reinforcement learning algorithms. Preliminary versions of the results in this paper appeared in the conference paper [29], however, the estimates provided in this paper significantly improve upon the results in [29]. In particular, we also address the case in which the sensitivity does not decay exponentially but only polynomially. Moreover, we comment on the relation of our result to known approximation-theoretic methods, such as anchored decompositions.
Furthermore, we establish connections between the existence of separable approximations and neural networks, illustrating how structured architectures can benefit separability to achieve scalable representations of high-dimensional functions. Through theoretical results and numerical experiments, we demonstrate the potential of this approach in reducing computational complexity while maintaining high approximation accuracy.
@@ 1.3 Notation
[…]
@@ 2 General setting for decaying sensitivity
In this section, we introduce sensitivity for graph-based problems through a general framework. Though our primary results in Section 4 address decaying sensitivity in optimal control problems, the underlying concepts have broader applicability. This general formulation is valuable because the existence of separable approximations - a key finding related to decaying sensitivity - extends beyond optimal control to other graph-based problems.
@@
Let us consider a system $s\in\mathbb{N}$ agents with states $z_{1},\dots,z_{s}$ with $z_{j}\in\mathbb{R}^{n_{j}}$ for $j=1,\dots,s$, which together constitute the state space $\mathbb{R}^{n}$, where $n=\sum_{j=1}^{s}n_{j}$. Denote by $x=(z_{1},\dots,z_{s})\in\mathbb{R}^{n}$ the combined state of all agents. The interactions among the agents are represented by a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, consisting of $s$ vertices $\mathcal{V}=\{1,\dots,s\}$, corresponding to the agents, with an edge $(i,j)\in\mathcal{E}$ indicating that agent $z_{i}$ exerts some influence on $z_{j}$. For $i,j\in\{1,\dots,s\}$, define $\text{dist}_{\mathcal{G}}(i,j)$ to be the length of the shortest path from node $i$ to node $j$ in $\mathcal{G}$ (possibly $\infty$ if no path exists and 00 if $i=j$). We are interested in computing a quantity of interest given by a function $V\colon\Omega\subset\mathbb{R}^{n}\to\mathbb{R}$.
Note that we use a directed graph $\mathcal{G}$ to describe the interaction among agents. However, in the special case of symmetric influence, these interactions can be represented using an undirected graph, which can be viewed as a particular case of a directed graph. While the discussion in Section 4 and the presented examples focus on the symmetric case, leading to undirected graphs, we adopt the more general formulation with directed graphs, as the main results in Section 3 hold for both cases. In the following, we formalize the concept of a decaying sensitivity between agents with an increasing graph distance. To this end, for each $j\in\{1,\dots,s\}$ define the mapping
@@
Each such mapping $g_{j}$ allows us to quantify the influence that the value of $z_{j}$ has on the evaluation of $V$ at some point $x$ in the state space compared to the anchor 00. While the mappings $g_{j}$ above could be defined using a general anchor in place of 00 for $z_{j}$ in the second term of the difference, we choose to compare with 00 to avoid notational overload. We are now interested in how much the value of such a function $g_{j}$ at some point in the state space is affected by a change of the values $z_{i}$ with $i\in\{1,\dots,s\},i\neq j$. We characterize this effect by examining the Lipschitz constant of the mapping $g_{j}$ with respect to $z_{i}$. This leads to the following definition:
@@
If the considered graph $\mathcal{G}$ and the set $\Omega$ are clear from the context, we refer to this property as $\gamma$-sensitivity. Intuitively, Definition 2.1 implies that the influence an agent $z_{i}$ has on $z_{j}$, as measured by the mapping $g_{j}$, decreases as the graph distance between $i$ and $j$ increases. Note that in the situation of Definition 2.1 we may have $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)=\infty$ for some $i,j\in\{1,\dots,s\}$ if there is no path from node $i$ to node $j$ in $\mathcal{G}$. In this case, (2) is to be understood as $L_{i,j}=0$, thus stating that $g_{j}$ is independent of $z_{i}$.
@@ 3 Separable approximation through decaying sensitivity
In this section, we first clarify the relationship between the decaying sensitivity property introduced in Definition 2.1 and the Hessian of the function $V$, under the assumption that $V$ is twice continuously differentiable. We then revisit the approximation method for $V$ proposed in [29] and apply Definition 2.1 to establish an enhanced error bound that is independent of the number of agents $s$, exhibiting either exponential or polynomial decay in sensitivity. Finally, we highlight a connection between our construction and the anchored decomposition, a well-established concept in approximation theory. Throughout this section, we work within the framework established in Setting 2.1.
@@ 3.1 Relation between decaying sensitivity and the Hessian
Assuming twice differentiability of the function $V$, we can deduce $\gamma$-sensitivity from a corresponding property of its Hessian, as established in the following result.
@@
then $V$ has $\gamma$-sensitivity on $\Omega$.
@@
Fix some $i,j\in\{1,\dots,s\}$, $i\neq j$. Let $z_{-j}=(z_{1},\dots,z_{j-1},z_{j+1},\dots,z_{s})\in\mathbb{R}^{n-n_{j}}$. By assumption, the mapping
@@
$\left\lVert{\frac{\partial}{\partial z_{i}}g_{j}(z_{j},z_{-j})}\right\rVert\leq\gamma(\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j))\left\lVert{z_{j}}\right\rVert,$
whence $L_{i,j}\coloneqq\gamma(\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j))\left\lVert{z_{j}}\right\rVert$ is a Lipschitz constant of $g_{j}$ with respect to $z_{i}$ at the point $(z_{1},\dots,z_{s})$. Then, applying Definition 2.1, the result follows. ∎
@@
Then $V$ admits a $\gamma$ decaying sensitivity property. In particular, if $P$ is $(C,\rho)$ SED for some $C>0$ and $\rho\in(0,1)$, then $V$ admits an exponential decaying sensitivity given by $\gamma(\cdot)=C\rho^{(\cdot)}$.
@@ 3.2 Construction of a Separable Approximation
Let $n,d\in\mathbb{N}$ with $d\leq n$. A function $F\colon\mathbb{R}^{n}\to\mathbb{R}$ is called $d$-separable if $F$ can be expressed as $F=\sum_{j=1}^{s}F_{j}$ for some $s\leq n$ and functions $F_{j}\colon\mathbb{R}^{d_{j}}\to\mathbb{R}$ with $d_{j}\leq d$ for all $1\leq j\leq s$.
Through a systematic elimination of the subsystems, a separable approximation of $V$ was constructed in [29]. The approximation depends on a variable $l\leq s$ and can be expressed as follows:
@@
where $z_{j,l}$ represents the vector containing the states of subsystem $z_{j}$ along with all subsystems $z_{i}$ with $i\geq j$ and $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)\leq l$. Let $d_{j,l}\in\mathbb{N}$ denote the dimension associated with $z_{j,l}$, i.e., $d_{j,l}$ is the sum of the dimensions of the subsystems that have an index smaller than $j$ and a graph distance of at most $l$ from $z_{j}$. Furthermore, let $x_{j,l}\in\mathbb{R}^{n}$ denote the lifting of $z_{j,l}$ into $\mathbb{R}^{n}$ by inserting the value 00 for all subsystems not included in $z_{j,l}$. The component functions $\Psi_{l}^{j}$ in (4) are defined as
@@
where $\Lambda_{j}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ sets the values of subsystem $j$ to 00, i.e., for $x=(z_{1},\dots,z_{s})$ it is
@@
Consequently, the expression in (4) constitutes a $d$-separable approximation with $d\coloneqq\max_{j\in\{1,\dots,s\}}d_{j,l}$. Figure 1 illustrates a graph composed of five subsystems, where the graph distance is given by $\text{dist}_{\mathcal{G}}(i,j)=|i-j|$. According to our notation, this example corresponds to $z_{1,2}=(z_{1},z_{2},z_{3})$, $x_{1,2}=(z_{1},z_{2},z_{3},0,0)$, and $\Psi_{2}^{1}(z_{1,2})=V(z_{1},z_{2},z_{3},0,0)-V(0,z_{2},z_{3},0,0)$. Note that the approximation (4) is based on a step-by-step elimination of the subsystems, beginning with $z_{1}$, followed by $z_{2}$, and so on, see Figure 2 in [29] for an illustration. Thus, for graphs where not all nodes are structurally identical, the indexing of the subsystems influences the sizes of the graph-neighborhoods and thus the order of separability.
Figure 1: Graph neighborhoods $z_{1,2}$ (left) and $z_{2,2}$ (right) in a sequential graph.
In the following, we establish an error bound based on the assumption of the decaying sensitivity property introduced in Definition 2.1. In contrast to Theorem 4 in [29], this error bound is designed to be independent of both the number of agents $s$ and the dimension $n$. However, as a trade-off, the result is restricted to $\mathcal{L}_{2}$ balls, instead of $\mathcal{L}_{\infty}$ balls, as presented in [29, Theorem 4]. This modification is achieved through the revised concept of decaying sensitivity outlined in Definition 2.1, as well as by leveraging the interplay between the type of decay and the growth of the graph as the dimension increases.
We begin by establishing an error bound for the approximation presented in (4). This bound is expressed in terms of the Euclidean norm of the state $x\in\mathbb{R}^{n}$ and the norm of a matrix $D_{l}\in\mathbb{R}^{s\times s}$, which reflects the decaying sensitivity property.
@@
Assume that $V$ has the $\gamma$-decaying sensitivity property. Then, for all $l\leq s$ and $x\in\Omega$, we have
@@
Initially, we observe that for any $i,j\in\{1,\dots,s\}$ we have that either $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)\leq s-1$ or $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)=\infty$. In the latter case it follows that $V(x)-V(\Pi_{j}x)$ is independent of $z_{i}$. Thus, by considering graph-neighborhoods of the size $s-1$ we can express $V(x)$ as $V(x)=V(0)+\sum_{j=1}^{s}\Psi_{s-1}^{j}(z_{j,s-1})$. By rewriting this expression as a telescoping sum we get
$\begin{split}&\Big{|}{V(x)-\sum_{j=1}^{s}\Psi_{l}^{j}(z_{j,l})-V(0)}\Bigg{|}=\Bigg{|}{\sum_{j=1}^{s}\Psi_{s-1}^{j}(z_{j,s-1})-\sum_{j=1}^{s}\Psi_{l}^{j}(z_{j,l})}\Bigg{|}\\ \leq&\sum_{j=1}^{s}\sum_{k=l}^{s-2}\left\lvert{\Psi_{k+1}^{j}(z_{j,k+1})-\Psi_{k}^{j}(z_{j,k})}\right\rvert.\end{split}$ (7)
Now, fix some $j\in\{1,\dots,s\}$ and let $\mathcal{G}_{j,l+1}=\{i_{1},\dots,i_{\tau}\}$. Then, by the definitions of $\Psi_{l}^{j}$ in (5) and $g_{j}$ in (1) we have
$\Psi_{l}^{j}(z_{j,l})-\Psi_{l+1}^{j}(z_{j,l+1})=V(x_{j,l})-V(\Lambda_{j}x_{j,l})-V(x_{j,l+1})+V(\Lambda_{j}x_{j,l+1})$
$=$ $\sum_{k=1}^{\tau}g_{j}(x_{j,l+1}^{(k)})-g_{j}(x_{j,l}^{(k)}),$
where $x_{j,l+1}^{(k)}$ and $x_{j,l}^{(k)}$ are obtained by setting the entries corresponding to $z_{i_{1}},\dots,z_{i_{k}}$ to 00 in the vectors $x_{j,l+1}$ and $x_{j,l}$, respectively. By successively applying the $\gamma$ decaying sensitivity property as defined in Definition 2.1, we obtain
$\left\lvert{\Psi_{l+1}^{j}(z_{j,l+1})-\Psi_{l}^{j}(z_{j,l})}\right\rvert\leq\gamma(l+1)\sum_{k=1}^{\tau}\left\lVert{z_{i_{k}}}\right\rVert\left\lVert{z_{j}}\right\rVert.$
@@
In the following, we state the error bound derived from Lemma 3.2 for both exponential and polynomial decay scenarios. For this, it is essential to quantify the growth of neighborhoods within the underlying graph. Denote for $i\in\{1,\dots,s\}$ and $l\in\mathbb{N}$ the set $\mathcal{G}_{i,l}=\{k\in\{1,\dots,s\}\mid\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,k)=l\}$. We define a function $r\colon\mathbb{N}\to\mathbb{N}$ as a growth bound for the graph $\mathcal{G}$ if it satisfies the condition:
$\forall l\in\mathbb{N},j\in\{1,\dots,s\}:\;\left\lvert\mathcal{G}_{j,l}\right\rvert\leq r(l).$ (8)

### v2

@@ 1 Introduction
One promising approach to mitigating the curse of dimensionality lies in exploiting separability, that is, the ability to approximate a global function as a sum of local contributions involving only a subset of variables. Separability properties can be intrinsic to the problem under study, as in systems where interactions decay with distance or graph connectivity, such as in physical systems governed by locality principles or networked systems with sparse interactions. At a computational level, separability can be enforced by a suitable approximation ansatz, including separation of variables, ANOVA[26], and most recent tools such as tensor decompositions[6, 25]. Theoretical and computational frameworks for separable approximations offer significant reductions in complexity while retaining essential structural features of the underlying functions.
@@
The concept of decaying sensitivity is intrinsically linked to the off-diagonal decay property, a topic extensively explored in the field of numerical linear algebra. A foundational theoretical framework for understanding this behavior is introduced in [4], where decay bounds are derived for the entries of general matrix functions. This analysis is further advanced in [15], which examines decay properties in the context of optimal control, specifically for solutions to Lyapunov equations. Additionally, related investigations are presented in [22], where the sparsity structure of solutions to Riccati equations is analyzed, providing complementary insights into the decay and sparsity phenomena in matrix equations. Our particular focus of application are optimal control problems, for which such decaying sensitivity properties in time and space have been intensively studied in the past years, see, e.g., [31, 24, 14] for decaying sensitivity in time and [30, 34, 27, 12] for decaying sensitivity in space for problems governed by discrete-time dynamics, ODEs and PDEs. See also [29] for a related property for static nonlinear optimization problems. The key prerequisite for the occurrence of decaying sensitivity in space is that the optimal control problem consists of subsystems that are interconnected by a graph structure. For instance in multi-agent control systems or robotic swarms, each agent’s state and control variables are influenced predominantly by neighboring agents, leading to a natural decaying sensitivity structure. In [9], it was shown that low-rank approximations to linear-quadratic value functions explicitly depend on the connectivity of the dynamics. By leveraging this decaying sensitivity property, we develop computationally efficient methods for approximating optimal value functions through separable structures.
@@ 1.1 Contribution
The main purpose of this paper is to show that decaying sensitivity in space enables the construction of accurate separable approximations. While our result is formulated for arbitrary functions and thus provides a general framework for separable approximations, we focus on the approximation of value functions arising in the context of dynamic programming or reinforcement learning algorithms. Preliminary versions of the results in this paper appeared in the conference paper [33], however, the estimates provided in this paper significantly improve upon the results in [33]. In particular, we also address the case in which the sensitivity does not decay exponentially but only polynomially. Moreover, we comment on the relation of our result to known approximation-theoretic methods, such as anchored decompositions.
Furthermore, we establish connections between the existence of separable approximations and neural networks, illustrating how structured architectures can benefit from separability to achieve scalable representations of high-dimensional functions. Through theoretical results and numerical experiments, we demonstrate the potential of this approach in reducing computational complexity while maintaining high approximation accuracy.
@@ 1.3 Notation
Throughout this paper, we adopt the following notation. For a vector $x\in\mathbb{R}^{n}$, we denote its one-dimensional components as $x_{j}$ for $1\leq j\leq n$. The entries of a given matrix $A$, which may include block-entries, are denoted by $A[i,j]$. The identity matrix in $\mathbb{R}^{n\times n}$ is denoted by $I_{n}$, and the null matrix of the same dimension is represented by $0_{n}$. Additionally, unless otherwise specified, the notation $\|\cdot\|$ refers to the Euclidean 2-norm. The comparison space $\mathcal{L}$ consists of all continuous and strictly decreasing functions $\gamma\colon\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ that satisfy the condition $\lim_{\tau\to\infty}\gamma(\tau)=0$. For a set $\Omega\subset\mathbb{R}^{n}$ and a function $f\colon\Omega\to\mathbb{R}^{m}$, we set $\lVert f\rVert_{\infty,\Omega}\coloneqq\sup_{x\in\Omega}\lVert f(x)\rVert.$ If $f$ is continuously differentiable, we further define
@@ 2 General setting for decaying sensitivity
In this section, we introduce sensitivity for graph-based problems through a general framework. Though our primary results in Section 4 address decaying sensitivity in optimal control problems, the underlying concepts have broader applicability. This general formulation is valuable because the existence of separable approximations – a key finding related to decaying sensitivity – extends beyond optimal control to other graph-based problems.
@@
Let us consider a system of $s\in\mathbb{N}$ agents with states $z_{1},\dots,z_{s}$ with $z_{j}\in\mathbb{R}^{n_{j}}$ for $j=1,\dots,s$, which together constitute a vector in the state space $\mathbb{R}^{n}$, where $n=\sum_{j=1}^{s}n_{j}$. Denote by $x=(z_{1},\dots,z_{s})\in\mathbb{R}^{n}$ the combined state of all agents. The interactions among the agents are represented by a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, consisting of $s$ vertices $\mathcal{V}=\{1,\dots,s\}$, corresponding to the agents, with an edge $(i,j)\in\mathcal{E}$ indicating that agent $z_{i}$ exerts some influence on $z_{j}$. For $i,j\in\{1,\dots,s\}$, define $\text{dist}_{\mathcal{G}}(i,j)$ to be the length of the shortest path from node $i$ to node $j$ in $\mathcal{G}$ (possibly $\infty$ if no path exists and 0 if $i=j$). We are interested in computing a quantity of interest given by a function $V\colon\Omega\subset\mathbb{R}^{n}\to\mathbb{R}$.
Note that we use a directed graph $\mathcal{G}$ to describe the interaction among agents. However, in the special case of symmetric influence, these interactions can be represented using an undirected graph, which can be viewed as a particular case of a directed graph. While the discussion in Section 4 and the presented examples focus on the symmetric case, leading to undirected graphs, we adopt the more general formulation with directed graphs, as the main results in Section 3 hold for both cases. We emphasize that the edges in our graph are defined as tuples of nodes without assigned weights and the length of a path is determined by the number of its edges. An interesting extension would be to consider weighted graphs, where the weight of an edge represents the strength of the coupling between the nodes and the graph distance between two nodes is defined as the sum of the weights along a path. A detailed analysis of this setting is, however, deferred to future research.
In the following, we formalize the concept of a decaying sensitivity between agents with an increasing graph distance. To this end, for each $j\in\{1,\dots,s\}$ define the mapping
@@
Each such mapping $g_{j}$ allows us to quantify the influence that the value of $z_{j}$ has on the evaluation of $V$ at some point $x$ in the state space compared to the anchor 0. While the mappings $g_{j}$ above could be defined using a general anchor in place of 0 in the second term of the difference, we choose to compare with 0 to avoid notational overload. We are now interested in how much the value of such a function $g_{j}$ at some point in the state space is affected by a change of the values $z_{i}$ with $i\in\{1,\dots,s\},i\neq j$. We characterize this effect by examining the Lipschitz constant of the mapping $g_{j}$ with respect to $z_{i}$. We call $g_{j}$ Lipschitz with respect to $z_{i}$ at a point $x=(z_{1},\dots,z_{s})\in\Omega$ with Lipschitz constant $L>0$ if for all $\tilde{x}=(z_{1},\dots,z_{i-1},\tilde{z}_{i},z_{i+1},\dots,z_{s})\in\Omega$ it holds that
$|g_{j}(x)-g_{j}(\tilde{x})|\leq L\|z_{i}-\tilde{z}_{i}\|.$
@@
If the considered graph $\mathcal{G}$ and the set $\Omega$ are clear from the context, we refer to this property as $\gamma$-decaying sensitivity. Intuitively, Definition 2.1 implies that the influence an agent $z_{i}$ has on $z_{j}$, as measured by the mapping $g_{j}$, decreases as the graph distance between $i$ and $j$ increases. Note that in the situation of Definition 2.1 we may have $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)=\infty$ for some $i,j\in\{1,\dots,s\}$ if there is no path from node $i$ to node $j$ in $\mathcal{G}$. In this case, (2) is to be understood as $L_{i,j}=0$, thus stating that $g_{j}$ is independent of $z_{i}$.
@@ 3 Separable approximation through decaying sensitivity
In this section, we first clarify the relationship between the decaying sensitivity property introduced in Definition 2.1 and the Hessian of the function $V$, under the assumption that $V$ is twice continuously differentiable. We then revisit the approximation method for $V$ proposed in [33] and apply Definition 2.1 to establish an enhanced error bound that is independent of the number of agents $s$ for settings exhibiting either exponential or polynomial decay in sensitivity. Finally, we highlight a connection between our construction and the anchored decomposition, a well-established concept in approximation theory. Throughout this section, we work within the framework established in Setting 2.1.
@@ 3.1 Relation between decaying sensitivity and the Hessian
Assuming twice differentiability of the function $V$, we can deduce $\gamma$-decaying sensitivity from a corresponding property of its Hessian, as established in the following result.
@@
then $V$ has $\gamma$-decaying sensitivity on $\Omega$.
@@
Fix some $\in\{1,\dots,s\}, iji\neq j$. Let $z_{-j}=(z_{1},\dots,z_{j-1},z_{j+1},\dots,z_{s})\in\mathbb{R}^{n-n_{j}}$. By assumption, the mapping
@@
$\left\lVert{\frac{\partial}{\partial z_{i}}g_{j}(z_{1},\dots,z_{s})}\right\rVert\leq\gamma(\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j))\left\lVert{z_{j}}\right\rVert.$
Thus, $L_{i,j}\coloneqq\gamma(\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j))\left\lVert{z_{j}}\right\rVert$ is a Lipschitz constant of $g_{j}$ with respect to $z_{i}$ at the point $x=(z_{1},\dots,z_{s})$. Then, applying Definition 2.1, the result follows. ∎
@@
Then $V$ admits a $\gamma$-decaying sensitivity property. In particular, if $P$ is $(C,\rho)$ SED for some $C>0$ and $\rho\in(0,1)$, then $V$ admits an exponential decaying sensitivity given by $\gamma(\cdot)=C\rho^{(\cdot)}$.
@@ 3.2 Construction of a Separable Approximation
Let $n,d\in\mathbb{N}$ with $d\leq n$ and $\Omega\subset\mathbb{R}^{n}$. A function $F\colon\mathbb{R}^{n}\to\mathbb{R}$ is called $d$-separable if there exist $k\in\mathbb{N}$ and functions $F_{j}\colon\mathbb{R}^{d_{j}}\to\mathbb{R}$ with $d_{j}\leq d$ for all $1\leq j\leq k$ such that for all $x\in\Omega$
$F(x)=\sum_{j=1}^{k}F_{j}(y_{j}),$
where for all $j\in\{1,\dots,k\}$ it holds that $y_{j}=(x_{\kappa_{j,1}},\dots,x_{\kappa_{j,d_{j}}})$ for some $\kappa_{j,i}\in\{1,\dots,n\}$, $1\leq i\leq d_{j}$.
We consider the following construction for approximating $V$ by a separable function, first presented in [33]. Although the construction therein was developed in the context of an optimal value function $V$, it directly extends to Setting 2.1, that is, to a general function $V$ defined on substates with an associated graph $\mathcal{G}$. The approximation depends on a parameter $l$, which determines the size of local neighborhoods in the graph $\mathcal{G}$ used for the separable representation. For any fixed $l\in\mathbb{N}$ it can be expressed as follows:
@@
where $z_{j,l}$ represents the vector containing the states of subsystem $z_{j}$ along with all subsystems $z_{i}$ with $i\geq j$ and $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)\leq l$. Let $d_{j,l}\in\mathbb{N}$ denote the dimension associated with $z_{j,l}$, i.e., $d_{j,l}$ is the sum of the dimensions of the subsystems that have an index larger than $j$ and a graph distance of at most $l$ from $z_{j}$. Furthermore, let $x_{j,l}\in\mathbb{R}^{n}$ denote the lifting of $z_{j,l}$ into $\mathbb{R}^{n}$ by inserting the value 0 for all subsystems not included in $z_{j,l}$. The component functions $\Psi_{l}^{j}$ in (4) are defined as
@@
where $\Lambda_{j}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ sets the values of subsystem $j$ to 0, i.e., for $x=(z_{1},\dots,z_{s})$ it is
@@
Figure 2 illustrates a graph composed of five subsystems, where the graph distance is given by $\text{dist}_{\mathcal{G}}(i,j)=|i-j|$. According to our notation, in this example we have $z_{1,2}=(z_{1},z_{2},z_{3})$, $x_{1,2}=(z_{1},z_{2},z_{3},0,0)$, and $\Psi_{2}^{1}(z_{1,2})=V(z_{1},z_{2},z_{3},0,0)-V(0,z_{2},z_{3},0,0)$. Note that (4) is based on a successive reduction of substates through the functions $\Psi_{l}^{j}$; see Figure 2 for an illustration. Specifically, for $x=(z_{1},\dots,z_{s})\in\Omega$, we begin with the approximation
$V(x)\approx\Psi_{l}^{1}(z_{1,l})+V(0,z_{2},\dots,z_{s}).$
In the next step, we approximately have
$V(0,z_{2},\dots,z_{s})\approx\Psi_{l}^{2}(z_{2,l})+V(0,0,z_{3},\dots,z_{s}).$
This procedure is continued iteratively until all substates are taken into account. The remaining term corresponds to the value at the origin, which then yields (4). Note that in general the resulting approximation will depend on the numbering of the subsystems $z_{j}$.
Remark 3.1 (Role of the parameter $l$).
The expression in (4) constitutes a $d$-separable approximation of $V$ with $d\coloneqq\max_{j\in\{1,\dots,s\}}d_{j,l}$. The parameter $l$ plays a crucial role in this construction. On one hand, it determines the dimension of the subspaces associated with the vectors $z_{j,l}$: smaller values of $l$ reduce the number of substates included in $z_{j,l}$, thereby decreasing $d_{j,l}$ and consequently lowering the order of separability. This reduction is advantageous, as it results in lower-dimensional functions in the separable approximation of $V$. On the other hand, larger values of $l$ account for longer graph distances, which generally lead to better approximation quality, see Lemma 3.2 in the following. In conclusion, $l$ provides a trade-off between sparsity and approximation quality. Moreover, note that while the presented construction is valid for all $l\in\mathbb{N}$, choosing $l\geq s$ yields the same graph neighborhoods as l=s1l=s-1, since the graph distance between any two nodes in a graph with $s$ nodes is at most s1s-1.
Figure 1: Graph neighborhoods $z_{1,2}$ (left ellipse) and $z_{2,2}$ (right ellipse) in a sequential graph.
Figure 2: Stepwise isolation of the substates $z_{j}$ through the functions $\Psi_{l}^{j}$ illustrated for $s=3$.
We now establish an error bound for the approximation in (4) based on the decaying sensitivity property introduced in Definition 2.1. In contrast to Theorem 4 in [33], this error bound is designed to be independent of both the number of agents $s$ and the dimension $n$. However, as a trade-off, the result is restricted to $\mathcal{L}_{2}$ balls, instead of $\mathcal{L}_{\infty}$ balls, as presented in [33, Theorem 4]. This modification is achieved through the revised concept of decaying sensitivity outlined in Definition 2.1, as well as by leveraging the interplay between the type of decay and the growth of the graph as the dimension increases.
@@
Assume that $V$ has the $\gamma$-decaying sensitivity property. Then, for any $l\in\mathbb{N}$ and $x\in\Omega$, we have
@@
Initially, we observe that for any $i,j\in\{1,\dots,s\}$ we have that either $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)\leq s-1$ or $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)=\infty$. In the latter case it follows that $V(x)-V(\Lambda_{j}x)$ is independent of $z_{i}$. Thus, by considering graph-neighborhoods of the size s1s-1 we can express $V(x)$ as $V(x)=V(0)+\sum_{j=1}^{s}\Psi_{s-1}^{j}(z_{j,s-1})$. This shows the claim for $l\geq s-1$. For the case $l\leq s-2$ we use a telescoping sum to write
[…]
Now, fix some $j\in\{1,\dots,s\}$ and $\tilde{l}\in\{l,\dots,s-2\}$. Let $\tau\in\mathbb{N}$ and $i_{1},\dots,i_{\tau}\in\{1,\dots,s\}$ such that $\{i\in\{1,\dots,s\}\mid\operatorname{\operatorname{dist_{\mathcal{G}}}}(j,i)=\tilde{l}\}=\{i_{1},\dots,i_{\tau}\}$. Then, by the definitions of $\Psi_{l}^{j}$ in (5) and $g_{j}$ in (1) we have
$\Psi_{\tilde{l}}^{j}(z_{j,{\tilde{l}}})-\Psi_{{\tilde{l}}+1}^{j}(z_{j,{\tilde{l}}+1})=V(x_{j,{\tilde{l}}})-V(\Lambda_{j}x_{j,{\tilde{l}}})-V(x_{j,{\tilde{l}}+1})+V(\Lambda_{j}x_{j,{\tilde{l}}+1})$
$=$ $\sum_{p=1}^{\tau}g_{j}(x_{j,{\tilde{l}}+1}^{(p)})-g_{j}(x_{j,{\tilde{l}}}^{(p)}),$
where for each $1\leq p\leq\tau$ the vectors $x_{j,{\tilde{l}}+1}^{(p)}$ and $x_{j,{\tilde{l}}}^{(p)}$ are obtained by setting all entries corresponding to $z_{i_{1}},\dots,z_{i_{p}}$ to 0 in the vectors $x_{j,{\tilde{l}}+1}$ and $x_{j,{\tilde{l}}}$, respectively. By successively applying the $\gamma$-decaying sensitivity property as defined in Definition 2.1, we obtain
$\left\lvert{\Psi_{\tilde{l}+1}^{j}(z_{j,\tilde{l}+1})-\Psi_{\tilde{l}}^{j}(z_{j,\tilde{l}})}\right\rvert\leq\gamma(\tilde{l}+1)\sum_{p=1}^{\tau}\left\lVert{z_{i_{p}}}\right\rVert\left\lVert{z_{j}}\right\rVert.$
@@
In particular, Lemma 3.2 shows that increasing the graph distance $l$ improves the approximation quality. Specifically, for $l\geq s-1$, we have $D_{l}=0$, which implies $\Psi_{l}+V(0)=V$. However, in this case, the local state $z_{1,l}$ already contains all substates connected to $z_{1}$, resulting in a high order of separability. In view of our objective to approximate $V$ using low-dimensional functions, it is therefore not desirable to choose $l\geq s-1$. Instead, $l$ should be selected as a trade-off between approximation accuracy and numerical efficiency, cf. Remark 3.1.
In the following, we specify the error bound derived from Lemma 3.2 for both exponential and polynomial decay scenarios. For this, it is essential to quantify the growth of neighborhoods within the underlying graph. We define a function $r\colon\mathbb{N}\to\mathbb{N}$ as a growth bound for the graph $\mathcal{G}$ if it satisfies the condition:
$\forall l\in\{1,\dots,s\},j\in\{1,\dots,s\}:\;\left\lvert\{i\in\{1,\dots,s\}\mid\operatorname{\operatorname{dist_{\mathcal{G}}}}(j,i)=l\}\right\rvert\leq r(l).$ (8)